# How Bright in Gravitational Waves are Millisecond Pulsars for the Galactic Center GeV Gamma-Ray Excess? A Systematic Study

### v1

@@ Abstract
The existence of dark matter (DM) is supported by various macroscopic observations, but its microscopic nature remains elusive. The Galactic Center GeV gamma-ray excess (GCE) has been a leading candidate signal for particle dark matter annihilation. However, an unresolved population of millisecond pulsars (MSPs) in the bulge provides the alternative explanation for the excess. Identifying these MSPs in electromagnetic bands is difficult due to source confusion, pulse broadening, and extinction. Gravitational waves (GWs) provide a complementary probe: a steadily rotating, non-axisymmetric MSP emits a nearly monochromatic GW signal in the sensitive band of ground-based detectors, with amplitude set by its ellipticity. In this work, we systematically investigate the GW emission from the MSP population proposed to explain the GCE and its detectability with current and future detectors. We consider three major scenarios for the origin of ellipticity and model the population properties of these MSPs. We also consider both isolated MSPs and MSPs in binary systems, as well as Doppler effects in the detection. We find that while the signal is below the reach of current interferometers, next-generation detectors such as the Einstein Telescope (ET) and Cosmic Explorer (CE) can detect a fraction of those MSPs, offering a novel test of the MSP interpretation of the GCE. Future directed searches toward the Galactic Center with continued improvements in sensitivities will either uncover this long-sought MSP population or place stringent limits on their ellipticities and abundance, with important implications for both the astrophysical and dark-matter interpretations of the GCE.
@@ I Introduction
The existence of dark matter is supported by a broad set of macroscopic observations, including galaxy rotation curves Rubin et al. (1980); Bosma (1981); Corbelli and Salucci (2000), lensing in clusters (including the Bullet Cluster) Clowe et al. (2004, 2006), large-scale structure Eisenstein et al. (2005); Springel et al. (2006), and precision measurements of the cosmic microwave background Hinshaw et al. (2013); Aghanim et al. (2020). Despite this breadth of macroscopic evidence and an increasingly precise determination of the cosmic DM abundance, the microscopic nature of DM remains unknown: no experiment has yet identified the particle(s) that constitute DM or measured their interactions beyond gravity Jungman et al. (1996); Bertone and Hooper (2018); Cirelli et al. (2024). Worldwide experiments conducting direct, indirect, and collider searches continue to narrow viable parameter space across many candidate frameworks (e.g., WIMPs, axions, sterile neutrinos), but no conclusive signal has been established Bulbul et al. (2014); Ackermann et al. (2015a); Aaboud et al. (2017); Du et al. (2018); Billard et al. (2022); Meng et al. (2021); An et al. (2023); Guo et al. (2023); Aalbers et al. (2023); Aprile et al. (2023); Cao et al. (2024).
@@
The GCE has been a leading candidate signal for particle DM. It is characterized by a bump spectrum centered around 2 GeV with a several-GeV width and a roughly spherically symmetric morphology extending several degrees around the Galactic Center (GC). It was first identified and interpreted as a possible signature of dark matter annihilation in Ref. Goodenough and Hooper (2009), and its robustness against uncertainties in galactic diffuse foreground modeling was subsequently confirmed in multiple independent analyses Macias and Gordon (2014); Abazajian et al. (2014); Zhou et al. (2015); Calore et al. (2015a); Huang et al. (2016); Ackermann et al. (2017); Zhong et al. (2020); Cholis et al. (2022); Zhong and Cholis (2024); Song et al. (2024). A variety of particle-DM models have been invoked to explain the GCE Hooper and Goodenough (2011); Daylan et al. (2016); Calore et al. (2015b); Agrawal et al. (2015); Izaguirre et al. (2014); Leane and Slatyer (2019); Murgia (2020), most prominently weakly interacting massive particles (WIMPs) with masses of a few tens of GeV annihilating predominantly into $b\bar{b}$, which hadronize and decay to gamma rays.
@@
However, there is also a well-motivated astrophysical interpretation: an unresolved population of MSPs in the inner Galaxy can reproduce the observed GCE spectrum and morphology at current Fermi-LAT resolution Abazajian (2011); Yuan and Zhang (2014); Petrović et al. (2015); Lee et al. (2016); Bartels et al. (2016); Macias et al. (2018); Bartels et al. (2018); Gautam et al. (2022). These MSPs are mostly in the Galactic bulge, so hereafter we refer to them as “bulge MSPs” for simplicity. MSPs are rapidly rotating neutron stars (NSs) whose rotational kinetic energy can power radio, gamma rays, and other electromagnetic emissions. Young pulsars spin down on Myr timescales; subsequent accretion can recycle “dead” pulsars to millisecond periods Alpar et al. (1982). These recycled MSPs typically have lower surface magnetic fields, $B_{\rm surf}\sim 10^{8}$$10^{9}$ G, and their long lifetimes of order Gyr make the existence of a substantial bulge population astrophysically plausible Bhattacharya and van den Heuvel (1991); Lorimer (2008); Gnedin et al. (2014); Arca-Sedda and Capuzzo-Dolcetta (2014); Brandt and Kocsis (2015); Macquart and Kanekar (2015).
@@
GWs provide a complementary probe that is insensitive to dust, scattering, and the source-confusion problems in the crowded stellar fields around the GC. A steadily rotating, non-axisymmetric neutron star emits a continuous, nearly monochromatic GW signal at $f_{\rm GW}\simeq 2f_{\rm rot}$, typically in the $\simeq 200-2000$ Hz range for MSPs, which coincides with the most sensitive band of ground-based interferometers  [^1] In addition, the mHz band of the Laser Interferometer Space Antenna could also probe those MSPs in the binary systems; see Ref. Korol and Igoshev (2025) for a recent study. , including LIGO Aasi et al. (2015), Virgo Acernese et al. (2015), KAGRA Akutsu et al. (2019), and next-generation CE Evans et al. (2023) and ET Punturo et al. (2010). Although it is very difficult to resolve the large number of MSPs in the crowded bulge region using gamma rays and other electromagnetic wavelengths, GWs can potentially resolve individual MSPs in the frequency domain, thanks to a resolution of $\Delta f\sim 1/T_{\rm obs}$.
@@
Building upon the previous work Calore et al. (2019); Miller and Zhao (2023); Bartel and Profumo (2024), we present a refined analysis that incorporates a more physically motivated signal prediction and a detailed treatment of the detector response. This allows us to investigate the GW emission expected from the MSP population responsible for the GCE and assess its detectability with present and future ground-based GW interferometers  [^2] As a complementary avenue to ground‑based interferometers, a recent proposal suggests probing kHz‑band backgrounds using a laser-polarization “Gravitational Photon Polarization Twist” in space Capdevilla (2025). With optimistic baselines ($\sim 10^{6}$ km) and millisecond pulses, its reach could overlap the collective strains expected from a bulge MSP population Capdevilla (2025). . The GW emission arises from a finite ellipticity, and we consider major physically motivated scenarios that can generate such ellipticities: (1) internal magnetic-field-induced deformation, (2) long-lived crustal mountains from breaking strain, and (3) a general parametrization in terms of a fraction of the pulsar spin-down energy. Importantly, the third scenario also naturally yields a theoretical maximum, assuming all the spin-down power is converted to GW emission. We model the bulge MSPs by combining a fixed boxy bulge morphology, constrained by the GCE, with two distinct population frameworks, an Empirical Disk-Proxy Model and an Evolutionary Model, which respectively benchmark and self-consistently derive the necessary intrinsic MSP parameters and total population size. For detectability, we outline two search strategies—coherent and incoherent—that explicitly account for Doppler shifts from Earth’s motion and binary orbits. Our results demonstrate that, under conservative assumptions, the bulge MSPs can produce GW signals approaching the sensitivity of next-generation ground-based interferometers, such as CE and the ET Srivastava et al. (2022); Hild et al. (2008). For the theoretical maximum, a fraction of MSPs can reach the LIGO-Virgo-Kagra (LVK) O4 sensitivity Capote et al. (2025); Abac et al. (2025).
@@ III.2 Crustal Mountains from Breaking Strain
As a newly formed neutron star cools, its outer layers crystallize into a solid crust. This crust solidifies in an equilibrium shape corresponding to the star’s rapid initial rotation. As the NS spins down, the hydrostatic equilibrium shape gradually evolves, while the rigid crust resists the deformation. Stress therefore accumulates until the energy difference between the new equilibrium shape and the actual crust configuration becomes sufficiently large that the crust can no longer withstand it. At this point, the crust breaks down and adjusts to the new equilibrium. This scenario provides a natural mechanism for generating a non-axisymmetric ellipticity in the NS Ushomirsky et al. (2000); Owen (2005); Andersson et al. (2010); Johnson-McDaniel and Owen (2013).
@@ IV.1 Models for MSP Population Parameters
A statistically significant population of MSPs has yet to be individually resolved in the bulge. Therefore, inferring the properties of this putative population requires model-based approaches, for which we explore two primary, physically distinct frameworks. The first is the Empirical Disk-Proxy Model (Pop. I), which assumes the bulge MSPs share the same present-day properties as the well-characterized population in the Galactic disk, with fiducial distributions derived from the Australian Telescope National Facility (ATNF) pulsar database Manchester et al. (2005). The second framework is the Evolutionary Model (Pop. II), which is a theoretical approach informed by the work of Ploeg et al. Ploeg et al. (2020). It interprets the GCE as originating from an older MSP population native to the bulge, allowing for a self-consistent derivation of its evolved properties (e.g., longer periods, lower luminosities) from universal birth parameters and the region’s specific star formation history Crocker et al. (2016). While the Pop. II provides a direct fit to the observed GCE, its predicted number of MSP, discussed in Sec. IV.4, exhibits mild tension with a recent machine learning analysis of the GCE List et al. (2025), which favors a lager MSP population with $N_{\rm MSP}>10^{5}$.
@@
To ensure greater precision, the Shklovskii effect Shklovskii (1970) and the differential Galactic rotation Damour and Taylor (1991) should be considered in correcting the $\dot{P}$ (also consequently $\dot{f}_{\rm rot}$). The GW detectability, discussed in the subsequent sections, mostly depends on the high-ellipticity tail of the ellipticity distribution, which scales with $\dot{P}$ (Eqs. (8) and (13)). However, as demonstrated in Fig. 1 of Ref. Woan et al. (2018), such corrections are negligible for rather lager values of $\dot{P}$. Therefore, the influence of these effects on GW detectability is insignificant in our work.
@@
By sampling from the birth distributions of $P_{i}$ and $B_{\rm surf}$ and evolving each MSP over its sampled age, we generate the present-day period ($P$) distribution for the bulge population. With the distributions for the present-day period $P$ and the (time-independent) magnetic field $B_{\rm surf}$ established, we then induce the distribution for the period derivative, $\dot{P}$, using the same force-free magnetosphere relation as before:
@@ IV.3 Morphology
The morphology of the GCE provides crucial information about its origin. While early analyses suggested a spherically symmetric profile, such as a generalized Navarro-Frenk-White (gNFW) profile consistent with dark matter annihilation, compelling recent evidence indicates that the GCE’s spatial distribution may correlate with the stellar mass of the Galactic bulge Bartels et al. (2016); Lee et al. (2016); Macias et al. (2018); Bartels et al. (2018). This finding thus supports an astrophysical origin, such as a population of unresolved MSPs  [^4] See also Ref. Muru et al. (2025) for a recent study, which shows that realistic MW simulations yield a flattened, boxy DM distribution, making the GCE morphology from WIMP annihilation hard to distinguish from that of MSPs..
@@ V.2 Complications due to Doppler Shifts
Doppler shift can, in principle, be corrected by defining a new time variable $t^{\prime}=t+({\mathbf{x}(t)\cdot\hat{\mathbf{r}}})/{c}$ Maggiore (2007), where $\mathbf{x}(t)$ is the position of the detector and $\hat{\mathbf{r}}$ is the unit vector pointing toward the source. However, to correct the Doppler shift, the location of the MSPs needs to be known accurately, e.g.,
@@ V.3 Our Approach: Two Detection strategies
The first is the coherent strategy, which assumes that all Doppler shifts can be corrected. This is feasible for isolated MSPs in blind searches with sufficiently small pixel sizes (Eq. (28)), provided that adequate computational power is available. In this case, an individual MSP becomes detectable once its GW strain amplitude ($h_{0}$; Eq. (2)) exceeds the sensitivity threshold ($h_{0}^{\min}$; Eq. (23)) with ${\rm SNR}=5$ Abbott et al. (2021a); Abac et al. (2025) and $S_{n}$ for different detectors, i.e.,
@@
Therefore, in the incoherent strategy, the observables are no longer individual monochromatic signals, but rather, the collective excess power within each Doppler broadened frequency window.

### v2

# arXiv:2511.15793v2[astro-ph.HE] 02 Mar 2026

@@ Abstract
The existence of dark matter (DM) is supported by various macroscopic observations, but its microscopic nature remains elusive. The Galactic Center GeV gamma-ray excess (GCE) has been a leading candidate signal for particle DM annihilation. However, an unresolved population of millisecond pulsars (MSPs) in the bulge provides the alternative explanation for the excess. Identifying these MSPs in electromagnetic bands is difficult due to source confusion, pulse broadening, and extinction. Gravitational waves (GWs) provide a complementary probe: a steadily rotating, non-axisymmetric MSP emits a nearly monochromatic GW signal in the sensitive band of ground-based detectors, with amplitude set by its ellipticity. In this work, we systematically investigate the GW emission from the MSP population proposed to explain the GCE and its detectability with current and future detectors. We consider three major scenarios for the origin of ellipticity and model the population properties of these MSPs. We also consider both isolated MSPs and MSPs in binary systems, as well as Doppler effects in the detection. We find that while the signal is below the reach of current interferometers, next-generation detectors such as the Einstein Telescope (ET) and Cosmic Explorer (CE) may detect a fraction of those MSPs, offering a novel test of the MSP interpretation of the GCE. Future directed searches toward the Galactic Center with continued improvements in sensitivities will either uncover this long-sought MSP population or place stringent limits on their ellipticities and abundance, with important implications for both the astrophysical and DM interpretations of the GCE.
@@ I Introduction
The existence of DM is supported by a broad set of macroscopic observations, including galaxy rotation curves Rubin et al. (1980); Bosma (1981); Corbelli and Salucci (2000), lensing in clusters (including the Bullet Cluster) Clowe et al. (2004, 2006), large-scale structure Eisenstein et al. (2005); Springel et al. (2006), and precision measurements of the cosmic microwave background Hinshaw et al. (2013); Aghanim et al. (2020). Despite this breadth of macroscopic evidence and an increasingly precise determination of the cosmic DM abundance, the microscopic nature of DM remains unknown: no experiment has yet identified the particle(s) that constitute DM or measured their interactions beyond gravity Jungman et al. (1996); Bertone and Hooper (2018); Cirelli et al. (2024). Worldwide experiments conducting direct, indirect, and collider searches continue to narrow viable parameter space across many candidate frameworks (e.g., WIMPs, axions, sterile neutrinos), but no conclusive signal has been established Bulbul et al. (2014); Ackermann et al. (2015a); Aaboud et al. (2017); Du et al. (2018); Billard et al. (2022); Meng et al. (2021); An et al. (2023); Guo et al. (2023); Aalbers et al. (2023); Aprile et al. (2023); Cao et al. (2024).
@@
The GCE has been a leading candidate signal for particle DM. It is characterized by a bump spectrum centered around 2 GeV with a several-GeV width and a roughly spherically symmetric morphology extending several degrees around the Galactic Center (GC). It was first identified and interpreted as a possible signature of DM annihilation in Ref. Goodenough and Hooper (2009), and its robustness against uncertainties in galactic diffuse foreground modeling was subsequently confirmed in multiple independent analyses Macias and Gordon (2014); Abazajian et al. (2014); Zhou et al. (2015); Calore et al. (2015a); Huang et al. (2016); Ackermann et al. (2017); Zhong et al. (2020); Cholis et al. (2022); Zhong and Cholis (2024); Song et al. (2024). A variety of particle DM models have been invoked to explain the GCE Hooper and Goodenough (2011); Daylan et al. (2016); Calore et al. (2015b); Agrawal et al. (2015); Izaguirre et al. (2014); Leane and Slatyer (2019); Murgia (2020), most prominently weakly interacting massive particles (WIMPs) with masses of a few tens of GeV annihilating predominantly into $b\bar{b}$, which hadronize and decay to gamma rays.
@@
However, there is also a well-motivated astrophysical interpretation: an unresolved population of MSPs in the inner Galaxy can reproduce the observed GCE spectrum and morphology at current Fermi-LAT resolution Abazajian (2011); Yuan and Zhang (2014); Petrović et al. (2015); Lee et al. (2016); Bartels et al. (2016); Macias et al. (2018); Bartels et al. (2018); Gautam et al. (2022) [^1] We note that some earlier studies questioned the MSP interpretation, arguing that the spectra of known MSPs are too soft at sub-GeV energies and that population estimates inferred from low-mass X-ray binary counts suggest unresolved MSPs could contribute only a small fraction of the GCE intensity Hooper et al. (2013); Cholis et al. (2015). However, subsequent analyses have largely addressed these concerns. The sub-GeV spectrum of the GCE is now known to be heavily affected by systematic uncertainties in Galactic diffuse emission modeling Ackermann et al. (2017). Furthermore, recent binary population synthesis models have demonstrated that an MSP population formed via the accretion-induced collapse of oxygen-neon white dwarfs can naturally reproduce the GCE’s intensity, spectrum, and morphology, while strictly obeying LMXB constraints Gautam et al. (2022).. These MSPs are mostly in the Galactic bulge, so hereafter we refer to them as “bulge MSPs” for simplicity. MSPs are rapidly rotating neutron stars (NSs) whose rotational kinetic energy can power radio, gamma rays, and other electromagnetic emissions. Young pulsars spin down on Myr timescales; subsequent accretion can recycle “dead” pulsars to millisecond periods Alpar et al. (1982). These recycled MSPs typically have lower surface magnetic fields, $B_{\rm surf}\sim 10^{8}$$10^{9}$ G, and their long lifetimes of order Gyr make the existence of a substantial bulge population astrophysically plausible Bhattacharya and van den Heuvel (1991); Lorimer (2008); Gnedin et al. (2014); Arca-Sedda and Capuzzo-Dolcetta (2014); Brandt and Kocsis (2015); Macquart and Kanekar (2015).
@@
GWs provide a complementary probe that is insensitive to dust, scattering, and the source-confusion problems in the crowded stellar fields around the GC. A steadily rotating, non-axisymmetric neutron star emits a continuous, nearly monochromatic GW signal at $f_{\rm GW}\simeq 2f_{\rm rot}$, typically in the $\simeq 200-2000$ Hz range for MSPs, which coincides with the most sensitive band of ground-based interferometers  [^2] In addition, the mHz band of the Laser Interferometer Space Antenna could also probe those MSPs in the binary systems; see Ref. Korol and Igoshev (2025) for a recent study. , including LIGO Aasi et al. (2015), Virgo Acernese et al. (2015), KAGRA Akutsu et al. (2019), and next-generation CE Evans et al. (2023) and ET Punturo et al. (2010). Such GW emission is persistent and nearly monochromatic, in stark contrast to the transient, broadband signals from compact-binary coalescences. Although it is very difficult to resolve the large number of MSPs in the crowded bulge region using gamma rays and other electromagnetic wavelengths, GWs can potentially resolve individual MSPs in the frequency domain, thanks to a resolution of $\Delta f\sim 1/T_{\rm obs}$.
@@
Building upon the previous work Calore et al. (2019); Miller and Zhao (2023); Bartel and Profumo (2024), we present a refined analysis that incorporates a more physically motivated signal prediction and a detailed treatment of the detector response. This allows us to investigate the GW emission expected from the MSP population responsible for the GCE and assess its detectability with present and future ground-based GW interferometers  [^3] As a complementary avenue to ground‑based interferometers, a recent proposal suggests probing kHz‑band backgrounds using a laser-polarization “Gravitational Photon Polarization Twist” in space Capdevilla (2025). With optimistic baselines ($\sim 10^{6}$ km) and millisecond pulses, its reach could overlap the collective strains expected from a bulge MSP population Capdevilla (2025). . The GW emission arises from a finite ellipticity, and we consider major physically motivated scenarios that can generate such ellipticities: (1) internal magnetic-field-induced deformation, (2) long-lived crustal mountains from breaking strain, and (3) a general parametrization in terms of a fraction of the pulsar spin-down energy. Importantly, the third scenario also naturally yields a theoretical maximum, assuming all the spin-down power is converted to GW emission. We model the bulge MSPs by combining a fixed boxy bulge morphology, constrained by the GCE, with two distinct population frameworks, an Empirical Disk-Proxy Model and an Evolutionary Model, which respectively benchmark and self-consistently derive the necessary intrinsic MSP parameters and total population size. For detectability, we outline two search strategies—coherent and incoherent—that explicitly account for Doppler shifts from Earth’s motion and binary orbits. Our results demonstrate that, under conservative assumptions, the bulge MSPs can produce GW signals approaching the sensitivity of next-generation ground-based interferometers, such as CE and the ET Srivastava et al. (2022); Hild et al. (2008). For the theoretical maximum, a fraction of MSPs can reach the LIGO-Virgo-Kagra (LVK) O4 sensitivity Capote et al. (2025); Abac et al. (2025a).
@@ III.2 Crustal Mountains from Breaking Strain
During the recycling of an MSP, accretion can produce large-scale interior temperature asymmetries misaligned from the spin axis. These gradients drive temperature-sensitive electron captures in the deep crust that provide a mass quadrupole, often referred to as a mountain. This accretion-driven structure offers a natural mechanism for GW emission, limited primarily by the crust’s breaking strain Bildsten (1998); Ushomirsky et al. (2000); Owen (2005); Andersson et al. (2010); Johnson-McDaniel and Owen (2013).
@@ IV.1 Models for MSP Population Parameters
A statistically significant population of MSPs has yet to be individually resolved in the bulge. Therefore, inferring the properties of this putative population requires model-based approaches, for which we explore two primary, physically distinct frameworks. The first is the Empirical Disk-Proxy Model (Pop. I), which assumes the bulge MSPs share the same present-day properties as the well-characterized population in the Galactic disk, with fiducial distributions derived from the Australian Telescope National Facility (ATNF) pulsar database Manchester et al. (2005). The second framework is the Evolutionary Model (Pop. II), which is a theoretical approach informed by the work of Ploeg et al. Ploeg et al. (2020). It interprets the GCE as originating from an older MSP population native to the bulge, allowing for a self-consistent derivation of its evolved properties (e.g., longer periods, lower luminosities) from universal birth parameters and the region’s specific star formation history Crocker et al. (2016). While the Pop. II provides a direct fit to the observed GCE, its predicted number of MSP, discussed in Sec. IV.4, exhibits mild tension with a recent machine learning analysis List et al. (2025), which builds upon earlier studies List et al. (2020, 2021) and favors a larger MSP population of $N_{\rm MSP}>10^{5}$.
@@
To ensure greater precision, the Shklovskii effect Shklovskii (1970) and the differential Galactic rotation Damour and Taylor (1991) should be considered in correcting the $\dot{P}$ (also consequently $\dot{f}_{\rm rot}$). The GW detectability, discussed in the subsequent sections, mostly depends on the high-ellipticity tail of the ellipticity distribution, which scales with $\dot{P}$ (Eqs. (8) and (13)). However, as demonstrated in Fig. 1 of Ref. Woan et al. (2018), such corrections are negligible for rather larger values of $\dot{P}$. Therefore, the influence of these effects on GW detectability is insignificant in our work.
@@
By sampling from the birth distributions of $P_{i}$ and $B_{\rm surf}$ and evolving each MSP over its sampled age, we generate the present-day period ($P$) distribution for the bulge population  [^5] See Eq. (2.23) and the surrounding text in Ref. Ploeg et al. (2020).. With the distributions for the present-day period $P$ and the (time-independent) magnetic field $B_{\rm surf}$ established, we then induce the distribution for the period derivative, $\dot{P}$, using the same force-free magnetosphere relation as before:
@@ IV.3 Morphology
The morphology of the GCE provides crucial information about its origin. While early analyses suggested a spherically symmetric profile, such as a generalized Navarro-Frenk-White (gNFW) profile consistent with DM annihilation, compelling recent evidence indicates that the GCE’s spatial distribution may correlate with the stellar mass of the Galactic bulge Bartels et al. (2016); Lee et al. (2016); Macias et al. (2018); Bartels et al. (2018). This finding thus supports an astrophysical origin, such as a population of unresolved MSPs  [^6] See also Ref. Muru et al. (2025) for a recent study, which shows that realistic MW simulations yield a flattened, boxy DM distribution, making the GCE morphology from WIMP annihilation hard to distinguish from that of MSPs..
@@ V.2 Complications due to Doppler Shifts
Doppler shift can, in principle, be corrected by defining a new time variable $t^{\prime}=t+({\mathbf{x}(t)\cdot\hat{\mathbf{r}}})/{c}$  [^7] See Eq. (7.149) and the surrounding text in Ref. Maggiore (2007)., where $\mathbf{x}(t)$ is the position of the detector and $\hat{\mathbf{r}}$ is the unit vector pointing toward the source. However, to correct the Doppler shift, the location of the MSPs needs to be known accurately  [^8] See Eq. (7.151) and the surrounding text in Ref. Maggiore (2007)., e.g.,
@@ V.3 Our Approach: Two Detection strategies
The first is the coherent strategy, which assumes that all Doppler shifts can be corrected. This is feasible for isolated MSPs in blind searches with sufficiently small pixel sizes (Eq. (28)), provided that adequate computational power is available. In this case, an individual MSP becomes detectable once its GW strain amplitude ($h_{0}$; Eq. (2)) exceeds the sensitivity threshold ($h_{0}^{\min}$; Eq. (23)) with ${\rm SNR}=5$ Abbott et al. (2021a); Abac et al. (2025a) and $S_{n}$ for different detectors, i.e.,
@@
Therefore, in the incoherent strategy, the observables are no longer individual monochromatic signals, but rather, the collective excess power within each Doppler broadened frequency window  [^9] This is somewhat analogous to the stochastic GW background Allen and Romano (1999); Sharma and Harms (2020); Biscoveanu et al. (2020); Abbott et al. (2021b); Mukherjee et al. (2022); Zhou et al. (2022, 2023); Zhong et al. (2023, 2024, 2025), though the MSP signals mainly from the galactic center..